OpenStudy (carlyleukhardt):

Eighteen is 60% of what number?

OpenStudy (♪chibiterasu):

You can just divide

OpenStudy (denonakavro):

divide 18 by 60 and do times 100

OpenStudy (♪chibiterasu):

!00% = 1 so 60% = 0.6

OpenStudy (carlyleukhardt):

Thanks :D :D <3

OpenStudy (♪chibiterasu):

Please don't give out direct answers. http://openstudy.com/code-of-conduct

OpenStudy (mathstudent55):

In math, "of" usually means "times." "is" means =. Also, "percent" means "divided by 100" or "over 100". An unknown is a variable, such as x. We can translate your sentence into an equation. Then we solve the equation.

OpenStudy (mathstudent55):

Colors can be very useful in math. \(\sf \color{red}{Eighteen} ~\color{green}{is} ~\color{blue}{60\%} ~\color{purple}{of} ~\color{orange}{what ~number}?\)

OpenStudy (mathstudent55):

\(\sf \color{red}{Eighteen} ~\color{green}{is} ~\color{blue}{60\%} ~\color{purple}{of} ~\color{orange}{what ~number}?\) \( \color{red}{~~~~~~18} ~\color{green}{~~=} ~\color{blue}{\dfrac{60}{100}} ~\color{purple}{\times} ~\color{orange}{~~~~~~x}\) We have this equation: \(18 = \dfrac{60}{100}x\) \(18 = 0.6x\) Solve for x. Divide both sides by 0.6